2. Aneutral variant of feather color in snow geese, called "blue phase caused by an unknown genetic variant. The less common blue phase results from a single dominant gene while the common white phase is homozygous recessive. When choosing a mate, young birds will most often -select a mate that resembles their parents' coloring: if the birds were hatched by a mixed pair, they will mate with either color phase. plumage color, and label variants as ff Imagine you figured out which gene controls white/blue making a migration (white) and FF, Ff(both blue). You are able to sample from a large population 2 Ff, and 35 ff stop at Market Lake Wildlife Management Area (near Roberts), and find 5 FF, on the population. individuals. Test for Hw equilibrium and draw conclusions about forces acting
Answer #1

Total allelic sites = 2×number of individuals = 2×(5+2+35) = 84

Alleles of F = (2×FF) + Ff = 10 + 2 = 12; allelic frequency (p) = 12/84 = 0.143

Alleles of f = (2×ff) + Ff = 70 + 2 = 72; allelic frequency (q) = 72/84 = 0.857

Predicted genotype frequency of FF = p2 = (0.143)2 = 0.020; predicted number of FF individuals = 0.020×42 = 0.84 (lets say 1)

Predicted genotype frequency of ff = q2 = (0.857)2 = 0.734; predicted number of FF individuals = 0.734×42 = 30.8 (lets say 31)

Predicted number of Ff individuals = 42 - (1+31) = 10

Since the predicted frequencies do not meet the observed frequencies, the population is not in HW equilibrium. Also, since the two homozygous genotypes are increasing at the cost of the heterozygous genotype, it is an example of directional selection, the direction being away from the heterozygous genotype.